5.The Mendes family bought a new house 8 years ago for $129,000. The house is now worth $186,000. Assuming a steady rate of growth, what was the yearly rate of appreciation? Round your answer to the nearest tenth of a percent (1.2% etc)

Answer by stanbon(75887)

The Mendes family bought a new house 8 years ago for $129,000. The house is now worth $186,000. Assuming a steady rate of growth, what was the yearly rate of appreciation? Round your answer to the nearest tenth of a percent (1.2% etc)
======
A(t) = Ao*r^t
--------
186000 = 129000*r^8
------
r^8 = 1.44186
----
r = (1.44186)^(1/8)
----
r = 1.0468
----
Ans: Growth rate = 4.68%
------------------
Cheers,
Stan H.
--------------